# William W. Cohen1, Haohan Wang4, Jiaxin Huang2 1CMU, 2WUSTL, 3UW, 4UIUC

### v1

@@ 4.3 Positive Correlation: LVLMs’ Grid Parsing & Puzzle-Solving Performance
Model performance on crossword puzzles exhibits strong dependency on grid parsing capabilities, with systematic biases in word orientations. We evaluate grid parsing ability by prompting models to parse completed puzzle grids and associated clues (prompt details in Appendix B.4). Successful grid parsing involves: (1) identifying grid indexing numbers, (2) mapping numbers to clues to determine word orientation, and (3) extracting words with boundary recognition. The grid parsing ability reflects an LVLM’s ability to interpret both spatial and textual information.
@@
Table 3: WCR on Grid Parsing.
We reveal two significant patterns: (1) As shown in Table 3, LVLMs extract Across words more accurately than Down words, revealing OCR limitations. (2) Fig. 3 demonstrates a strong positive correlation between grid parsing and overall puzzle-solving performance (both measured by WCR). Notably, when enabling thinking mode, Claude-3-7-Sonnet exhibits lower grid parsing performance, which aligns with its lower puzzle-solving performance. We hypothesize that the additional token generation in thinking mode diverts attention from image processing. These findings underscore the importance of precise spatial-textual interpretation for future LVLM design. For full results, please refer to Appendix A.2.
4.4 Agentic Evaluation Setting: Interactive Mode for VLMs
@@
Interactive Mode requires step-by-step puzzle solving instead of solving the entire puzzle in one pass. Specifically, the implementation of CrossWordBench allows for updating grid images each time a model provides an answer. We maintain a four-round conversation history due to context limitations. We introduce Interactive Success Step (ISS) to quantify the number of correctly solved words before the model’s first error. Figure 4 shows the cumulative distribution of ISS for each model—most models fail at the initial solution step on most puzzles.
Unlike the visual-of-thoughts approach, which focuses on generating intermediate reasoning plots, we employ external functions to update the grid. This enables evaluating LVLMs in agentic settings, with error feedback to provide cues for the model to refine its responses.
@@ 4.5 Beyond English: Evaluations on Multilingual, Dictionary-based, and Adapted Data
We extend zero-shot CoT prompting evaluation to Chinese, dictionary-based, and benchmark-adapted word–clue pairs in CrossWordBench, and show results in Table 4.
Reasoning models outperform conventional ones across metrics, and the performance gap between open-weight and proprietary models is reduced. o3-mini maintains a high ICR across all three settings, consistent with main result observations. Furthermore, the performance gap between open-weight and proprietary reasoning models decreases, possibly due to the simpler nature of these tasks and differences in training data. Interestingly, when solving Chinese puzzles, we observe that QwQ-32B consistently reasons in Chinese.
Effectiveness of grid construction: puzzles constructed with CommonsenseQA training data remain challenging. Despite the CommonsenseQA having saturated performance on most LLMs, its training set continue to pose significant challenges for reasoning models. For example, o3-mini achieves the highest WCR of 0.812, yet it still falls short of perfection. In contrast, best evaluated non-reasoning LLMs such as GPT-4o and open-weight LVLMs Pixtral-Large-Instruct obtain WCR of 0.524 and 0.439, respectively. These results highlight the effectiveness of this grid construction approach in leveraging existing benchmark data.
@@ 5 Analysis
5.1 Crossing Letter Count: Impact on Puzzle Performance
@@
Average word accuracy on reasoning models increases with crossing letter count. As shown in Figure 5, the average WCR for each range of crossing letter counts, divided into three groups, across five reasoning LLMs and nine non-reasoning LLMs. We observe that reasoning models achieve higher accuracy as the number of letter intersections increases, whereas this trend is not observed for non-reasoning LLMs. This pattern aligns with our main observations, where reasoning models tend to exhibit higher ICRs—indicating that they benefit from effectively utilizing grid constraints for solution space reduction.
@@
Figure 6: Self-reflection improvements on 7x7 English puzzles. Top row: reasoning LLMs. Bottom row: non-reasoning LVLMs.
5.2 Self-Reflection: Minimal Effect on Puzzle Performance
Backtracking and verifying previous answers are essential for solving crossword puzzles. To evaluate the effect of self-reflection, we include a manually crafted follow-up query that asks models to revisit their previous zero-shot CoT responses (see Appendix B.4 for details). As shown in Figure 6, there is no observable performance improvement for either reasoning LLMs or non-reasoning LVLMs. This suggests that additional interactions with manual prompting alone are insufficient to enhance reasoning capabilities for puzzle solving.
@@ 5.3 Test-Time Scaling: Diminishing Returns on Puzzle Performance
Figure 7: o3-mini performance on 7x7 English puzzles across three distinct reasoning efforts.
To investigate the impact of test-time scaling on CrossWordBench, we use o3-mini on English 7x7 puzzles and control its reasoning effort, which in turn determines the number of reasoning tokens generated during inference. As shown in Figure 7, a medium effort yields a significant performance improvement compared to low effort across all three metrics, while further doubling the reasoning tokens does not result in a significant additional performance gain, indicating diminishing marginal returns.
@@ 5.4 Discussion: Verifiable Crossword for Future Multimodal RL Training
Table 2 shows that reasoning models such as Open-Reasoner-Zero-32B, which are primarily trained on mathematical problems, demonstrate limited generalization to CrossWordBench. This highlights the inadequacy of relying predominantly on math-based problems with verifiable answers for reinforcement learning. As multimodal reasoning evolves, a significant challenge remains: the lack of multimodal datasets with verifiable rewards for training. We contend that crossword puzzles, owing to their inherent verifiability and multimodal nature, offer a promising alternative, and our framework can serve as a foundation to address this gap. We leave exploration of multimodal RL training on crossword puzzles for future work.
@@ 6 Conclusion
Appendix A More Results and Analysis
A.1 Behavior Analysis
To better analyze the errors made by models in puzzle-solving, we define two metrics: global length error and local length error. The global length error metric compares the number of words produced by the model with those in the reference answer list, assessing whether the model supplies an answer for every clue in the puzzle. In contrast, the local length error metric compares the length of each individual word to its corresponding reference, thereby quantifying the model’s adherence to the grid constraints. Table 5 shows that even the best-performing reasoning models, such as Claude-3-7-Sonnet with thinking mode and DeepSeek R1, exhibit global length errors on two puzzles. Almost all models—with the exception of o3-mini, which demonstrates the highest ICR—commit a significant number of local errors. Moreover, we observe that all global length errors arise from models providing shorter answers, meaning that they fail to address some clues. Furthermore, most models tend to produce answers that are longer than the corresponding reference answers.
@@ A.1 Behavior Analysis
A.2 Grid Parsing Performance
@@
Table 6: WCR of Grid Parsing for all models.
A.3 Beyond English: Evaluations on Multilingual, Dictionary-based, and Adapted Data
@@
A.4 Test-Time Scaling: Diminishing Returns on Puzzle Performance
@@ A.5 Token Usage
((a)) Token Usage of VLMs.
@@
Figure 8: Token usage for LVLMs and LLMs on 7x7 and 14x14 English puzzles. Reasoning Models are highlighted in red.
A.6 Input Format Ablations
In our main evaluations, both the clues and the grid are embedded within a single image as inputs for LVLMs. In this section, we investigate the effect of isolating the clues and providing them to the LVLMs as separate text inputs, while the image contains only an empty grid. The prompt used for this experimental setting is shown in Figure 11 in Appendix B.
@@ A.6 Input Format Ablations
Appendix B Implementation Details
B.1 Evaluation Metrics
@@ B.2 Parsing Details
[…]
@@ B.3 Generation Configuration
Claude-3-7-Sonnet (Anthropic, 2024) 8192 8192 0.0
Claude-3-7-Sonnet (Thinking) (Anthropic, 2024) 64000 64000 1.0
@@
B.4 Prompts
Figures 10, 11, 12, 13, 14, 15, 17, and 18 present all the prompts employed in this study—comprising image prompts for LVLMs and text prompts for LLMs. We observe that Claude-3-7-Sonnet sometimes produce partial outputs and request confirmation to continue. To mitigate this issue, we incorporate an additional system prompt (see Figure. 18 for the system prompt); note that this modification applies only to the Claude-3-7-Sonnet.

### v2

@@ 4.3 Positive Correlation: LVLMs’ Grid Parsing & Puzzle-Solving Performance
Model performance on crossword puzzles exhibits a strong dependence on grid parsing capabilities, with systematic biases in word orientations. We evaluate grid parsing ability by prompting models to parse completed puzzle grids and associated clues (prompt details are in Appendix D.4). Successful grid parsing requires: (1) identifying grid indexing numbers, (2) mapping numbers to clues to determine word orientation, and (3) extracting words with boundary recognition. The grid parsing ability reflects an LVLM’s ability to interpret both spatial and textual information.
Table 3: WCR on Grid Parsing.
@@
We reveal two significant patterns: (1) As shown in Table 3, LVLMs extract Across words more accurately than Down words, highlighting OCR limitations. (2) Figure 3 demonstrates a strong positive correlation between grid parsing and overall puzzle-solving performance (both measured by WCR). Notably, when enabling thinking mode, Claude 3.7 Sonnet exhibits lower grid parsing performance, which aligns with its lower puzzle-solving performance. We hypothesize that the additional token generation in thinking mode diverts attention from image processing. These findings highlight the critical role of precise spatial-textual interpretation for future LVLM design. For full results, please refer to Appendix C.2.
4.4 Agentic Evaluation Setting: Interactive Mode for LVLMs
@@
Interactive Mode requires step-by-step puzzle solving rather than completing the entire puzzle in a single pass. Specifically, the implementation of the controllable generation framework allows for updating grid images each time a model provides an answer.
@@
We maintain a four-round conversation history due to context window limitations. We introduce Interactive Success Step (ISS) to quantify how many words a model correctly solves before making its first error. Figure 4 shows the cumulative distribution of ISS for each model—most models fail at the initial solution step on most puzzles.
Unlike the visual-of-thoughts approach, which focuses on generating intermediate reasoning plots, we employ external functions to update the grid. This establishes a foundation for evaluating LVLMs in agentic settings, where the word placement functions can serve as callable tools for the model and provide error feedback to guide the model in refining its responses—for example, by indicating whether a proposed word matches the expected length and whether its intersecting characters align with previously filled cells.
@@ 4.5 Beyond English: Evaluations on Multilingual, Dictionary-based, and Adapted Data
We extend zero-shot CoT prompting evaluation to Chinese, dictionary-based, and benchmark-adapted word–clue pairs in CrossWordBench, with results shown in Table 4.
Reasoning models outperform conventional ones across metrics, and the performance gap between open-weight and proprietary models is reduced. o3-mini maintains a high ICR across all three datasets, consistent with the main evaluation results. Furthermore, the performance gap between open-weight and proprietary reasoning models is reduced, possibly due to the simplicity of these tasks and differences in training data. Interestingly, when solving Chinese puzzles, we observe that QwQ-32B consistently reasons in Chinese.
Effectiveness of grid construction: puzzles constructed with CommonsenseQA training data remain challenging. Despite CommonsenseQA having saturated performance on most LLMs, its training set continues to pose significant challenges for reasoning models. For example, o3-mini achieves the highest WCR of 0.812, yet it still falls short of perfection. In contrast, the best-performing non-reasoning LLMs such as GPT-4o and open-weight LVLMs Pixtral-Large-Instruct obtain WCR of 0.524 and 0.439, respectively. These results highlight the effectiveness of our grid construction strategy in repurposing existing benchmark data.
@@ 5 Analysis
Table 5: Grid Format Ablations.
Models Array Markdown
Claude-3-7-Sonnet 0.482±0.015 0.760±0.018
GPT-4o-2024-11-20 0.410±0.014 0.398±0.024
Gemini 2.0 Flash 0.301±0.014 0.309±0.015
DeepSeek-V3 0.303±0.014 0.294±0.016
o3-mini 0.587±0.023 0.592±0.024
5.1 Grid Format Ablations: LLMs Robustness to Textual Grid Representations
LLMs exhibit robustness to variations in grid text representation. In the text-only evaluation setting, the empty crossword puzzle grid is represented as a 2D binary array, where 1 denotes a blocked cell and 0 denotes an unfilled cell. This array is then prepended to the prompt containing clues. To evaluate the impact of different formatting choices, we test an alternative markdown-style representation in this section, with · indicates unfilled cells and - indicates blocked cells. An example of the two formats is provided in Appendix D.5 and Figure 19.
Due to space constraints in the main text, we defer additional ablation studies on LVLM input formats and few-shot prompting to Appendix C.6 and C.8, respectively. We also demonstrate how controlling prefill ratio can be used to adjust task difficulty. We encourage readers to refer to these appendices for a comprehensive analysis and further discussion.
5.2 Crossing Letters: Reasoning LLMs Improve with More Intersections
@@
Average word accuracy on reasoning models increases with crossing letter count.
Table 6: Average WCR results by crossing letter count and density on 7x7 English puzzles.
Density Crossing Letter Reason. Non-Reason.
Low Low(1) 0.4341 0.2960
Med(2) 0.5275 0.3355
High(3+) 0.6165 0.2794
Medium Low(1) 0.4271 0.2754
Med(2) 0.5477 0.3089
High(3+) 0.6029 0.4193
High Low(1) 0.5288 0.3097
Med(2) 0.5151 0.3217
High(3+) 0.5928 0.3030
Figure 5 presents the average WCR for each range of crossing letter counts, divided into three groups, across five reasoning and nine non-reasoning LLMs. We observe that reasoning models exhibit increasing accuracy with a greater number of letter intersections, whereas this trend is not observed for non-reasoning LLMs. This observation aligns with our main experiment results, in which reasoning models tend to exhibit higher ICRs, suggesting that they benefit from effectively utilizing grid constraints for solution space reduction. However, we do not observe the same pattern on 14x14 puzzles—likely because puzzles with larger grids are substantially more difficult even for reasoning models. We provide detailed results in Appendix C.7.
To further isolate the effect of crossing-letter density within a constrained grid, we perform a quasi-ablation study on 7x7 puzzles by stratifying puzzles into low, medium, and high-density groups based on the average number of crossing letters per word, as shown in Table 6. Overall, within different densities, reasoning LLMs demonstrate improved WCR with a greater number of crossing letters, whereas non-reasoning LLMs show relatively stable performance.
5.3 Self-Reflection: Limited Impact on Crossword Puzzle Solving
@@
Figure 6: Self-reflection improvements on 7x7 English puzzles across metrics. Top: reasoning LLMs. Bottom: non-reasoning LVLMs.
Backtracking and verifying previously filled answers are essential components of effective crossword puzzle solving. To evaluate the effect of self-reflection, we include a manually crafted follow-up query that prompts models to revisit their previous zero-shot CoT responses (please see Appendix D.4 for details). As shown in Figure 6, we observe no measurable performance improvement for reasoning LLMs or non-reasoning LVLMs. This find suggests that additional interactions with manual prompting alone are insufficient to enhance reasoning capabilities for puzzle solving.
@@ 5.4 Test-Time Scaling: Diminishing Returns on Puzzle Performance
Figure 7: o3-mini performance on 7x7 English puzzles under three levels of reasoning efforts.
To examine the effect of test-time scaling on CrossWordBench, we evaluate o3-mini on 7x7 English puzzles by varying its reasoning effort, which controls the number of reasoning tokens generated during inference. As shown in Figure 7, increasing the effort from low to medium yields a substantial performance improvement across all three metrics. However, further doubling the reasoning tokens provides no significant additional gains, indicating diminishing returns.
@@ 5.5 Discussion: Verifiable Crossword for Future Multimodal RL Training
Table 2 shows that reasoning models such as Open-Reasoner-Zero-32B, which are primarily trained on mathematical problems, exhibit limited generalization to CrossWordBench. This highlights the limitations of relying predominantly on math-based tasks with verifiable answers for reinforcement learning in more complex reasoning settings. As multimodal reasoning advances, a significant challenge persists: the lack of multimodal environments with verifiable answers suitable for rule-based reinforcement learning. We propose crossword puzzles as a compelling alternative, owing to their unique combinations of verifiability, multimodal structure, and goal-oriented task design. Crossword puzzles are also suitable for multi-turn training, where each word fill can represent a discrete, observable decision.
Our framework supports this setting by providing word placement functions that dynamically update the puzzle grid in both text and image formats, allowing seamless integration with multimodal agents and enabling tool-use training. We leave the exploration of training multimodal agents on crossword puzzles with real-time function feedback for future work.
@@ 6 Conclusion
Appendix A Acknowledgment
We thank all reviewers for their feedback, which contributed to improving this work.
Appendix B Limitations
We propose CrossWordBench to evaluate the reasoning capabilities of both LLMs and LVLMs. Unlike prior work, CrossWordBench does not incorporate any human-authored puzzles from online sources. Although our generation process is carefully designed to preserve key structural properties—such as clue distribution, word length, and blocked cell ratios—as summarized in Table 1, and the generated puzzles exhibit difficulty levels comparable to those of human—created puzzles when evaluated using standard word-level metrics, manual verification of puzzle quality could be helpful. Although our framework is also capable of generating crossword puzzles for potential use in training, this paper focuses solely on evaluation and leaves the exploration of reinforcement learning for future work.
Appendix C More Results and Analysis
C.1 Behavior Analysis
To better analyze the errors made by models in puzzle-solving, we define two metrics: global length error and local length error. The global length error metric compares the number of words produced by the model with those in the reference answer list, assessing whether the model supplies an answer for every clue in the puzzle. In contrast, the local length error metric compares the length of each individual word to its corresponding reference, thereby quantifying the model’s adherence to the grid constraints. Table 7 shows that even the best-performing reasoning models, such as Claude 3.7 Sonnet with thinking mode and DeepSeek R1, exhibit global length errors on two puzzles. Almost all models—with the exception of o3-mini, which demonstrates the highest ICR—commit a significant number of local errors. Moreover, we observe that all global length errors arise from models providing fewer answers than required, i.e., failing to address some clues. In contrast, most local length errors are due to models generating answers that exceed the expected word length.
@@ C.1 Behavior Analysis
C.2 Complete Results on Grid Parsing Performance
@@
C.3 Complete Results on Extended Data
@@
C.4 More on Test-Time Scaling
@@ C.5 Token Usage
((a)) Token Usage of LVLMs.
@@
Figure 8: Token usage on 7x7 and 14x14 English puzzles. Reasoning models are in red.
C.6 Ablations: LVLM Input Format and Modality Effects
[…]
In our main evaluations, both the clues and the grid are embedded within a single image as inputs for LVLMs. In this section, we investigate the effect of isolating the clues and providing them to the LVLMs as separate text inputs, while the image contains only an empty grid. The prompt used for this experimental setting is shown in Figure 11 in Appendix D.
@@ C.6 Ablations: LVLM Input Format and Modality Effects
C.7 More on Crossing Letter Count
Table 11: Average WCR results by crossing letter count on 7x7 and 14x14 English puzzles.
Grid Crossing Letter Reason. Non-Reason.
7x7 Low 0.4539 0.2922
7x7 Med 0.5308 0.3230
7x7 High 0.6183 0.3259
14x14 Low 0.4115 0.3469
14x14 Med 0.4186 0.3514
14x14 High 0.2306 0.1438
Table 11 presents the results shown in Figure 5, along with additional results on 14x14 English puzzles. On these larger puzzles, even reasoning LLMs struggle to effectively leverage crossing-letter constraints. We hypothesize that this is due to the increased complexity of larger puzzles, which remains challenging for reasoning models.
C.8 More on Few-Shot Prompting and Prefill Ratio Control
In our main evaluation, we treat each puzzle as a whole rather than as a set of independent question–answer (QA) pairs, in order to preserve the structural constraints inherent to crossword solving. This setup is not well-suited to few-shot prompting, as providing a full puzzle solution as a demonstration example would be lengthy and potentially distracting for the model. Nevertheless, we experiment with one-shot prompting on 7x7 English puzzles, using a completed example generated by DeepSeek-R1 on the English_Simple subset as the demonstration. As shown in Figure 12(a), the performance difference is not significant.
As discussed in Section 5.1, our controllable generation framework supports difficulty adjustment via the prefill ratio. In Table 12(b), we compare performance on 7x7 English puzzles under two prefill conditions: 0% and 50%. To ensure a fair comparison, we prevent any word from being fully revealed. As expected, a higher prefill ratio leads to improved performance, as more letters are made available to the model. These results suggest that prefill ratio is a more effective method for puzzle difficulty control than few-shot prompting.
Table 12: WCR under different prompting and prefill ratio settings on 7x7 English puzzles.
((a)) WCR with 0-shot and 1-shot prompting.
Model 0-Shot 1-Shot
GPT-4o (Text) 0.410 0.394
Claude-Sonnet (Text) 0.482 0.519
Gemini-Flash (Text) 0.301 0.313
Llama-3.3-70B (Text) 0.303 0.269
DeepSeek-V3 (Text) 0.303 0.307
((b)) WCR with 0% and 50% prefill.
Model 0% 50%
GPT-4o (Text) 0.410 0.483
GPT-4o (Img) 0.348 0.383
Claude-Sonnet (Text) 0.482 0.604
Claude-Sonnet (Img) 0.479 0.568
Gemini-Flash (Text) 0.301 0.362
Gemini-Flash (Img) 0.277 0.346
o3-mini (Text) 0.587 0.864
Appendix D Implementation Details
D.1 Evaluation Metrics
@@ D.2 Parsing Details
2:  fields $\leftarrow\{\}$ $\triangleright$ Initialize an empty dictionary
@@
4:  for all $answer\in\texttt{reference\_answers}$ do
5:   key $\leftarrow answer[\texttt{"direction"}]$
@@ D.3 Generation Configuration
Claude 3.7 Sonnet (Anthropic, 2024) 8192 8192 0.0
Claude 3.7 Sonnet (Thinking) (Anthropic, 2024) 64000 64000 1.0
@@
D.4 Prompts
Figures 10, 11, 12, 13, 14, 15, 17, and 18 present all the prompts employed in this study—comprising image prompts for LVLMs and text prompts for LLMs. We observe that Claude 3.7 Sonnet sometimes produces partial outputs and requests confirmation to continue. To mitigate this issue, we incorporate an additional system prompt (see Figure 18 for the system prompt); note that this modification applies only to the Claude 3.7 Sonnet.
@@ D.4 Prompts
D.5 Textual Grid Representations
$\begin{bmatrix}0&0&0&1&0&0&0\\ 0&1&0&0&1&0&1\\ 0&0&0&1&0&0&0\\ 0&1&1&1&0&1&0\\ 0&0&0&0&0&1&0\\ 1&0&1&0&1&1&0\\ 1&0&1&0&0&0&1\end{bmatrix}$
((a)) 2D array format (ARC-style).
[…]
((b)) Markdown-style grid with symbols.
Figure 19: Two text representations of the puzzle grid: array (left) and markdown (right).